OpenStudy (anonymous):

What ordered pair satisfies the inverse of y=3x-2? a.(1,2) b.(0,2) c.(4,2) d.(4,1)

OpenStudy (chillout):

Isolate x: \[\frac{y+2}{3}=x\] Exchange x into y:\[\frac{x+2}{3}=y\]Test the given points.

OpenStudy (chillout):

So, did you find the answer?

OpenStudy (anonymous):

i'm still answering

OpenStudy (anonymous):

wait :)

OpenStudy (anonymous):

is it letter c? @ChillOut

OpenStudy (chillout):

Yes, it is.

OpenStudy (anonymous):

kindly explain to me ... why ? @ChillOut

OpenStudy (chillout):

I'll give you the step-by-step: y = 3x -2. We add 2 to both sides, then it becomes y+2 = 3x. Dividing both sides by 3, we get to y+2/3 = x. Now all that we have to do is switch x by y. Just change their places! x+2/3 = y. If we test the option 3, we have (4 + 2)/3 = 2 -> 6/3 = 2, 2=2. See?

OpenStudy (chillout):

I meant test option c

OpenStudy (anonymous):

nice ! i get it now thank u again